OpenStudy (anonymous):

Two airplanes leave Dallas at the same time and fly in opposite directions. One airplane travels 80 mph faster than the other. After 3 hours, they are 2940 miles apart. What is the rate of each plane?

OpenStudy (kropot72):

Let s = speed of slower airplane Then speed of faster airplane = s + 80 Combined speed = s + s + 80 = 2s + 80 mph Distance = speed * time 2940 = 3(2s + 80) Can you now solve for s?

OpenStudy (anonymous):

i figured it out before but thanks for the effort. While your here, for motion problems it distance divided by rate to find time right

OpenStudy (kropot72):

Yes :)

OpenStudy (anonymous):

thanks!

OpenStudy (kropot72):

You're welcome :)